# Learning convolution operators on compact abelian groups

### v2

@@ A.2. Probabilistic tools
Associated to the feature map $\Phi$, we introduce some useful random variables, which play a central role in the proofs. By (53) the map $\Phi$ is continuous from $L^{1}$ into $\mathcal{B}(\mathcal{H},L^{2})$, then $\Phi(X),\Phi(X)^{*}$ and $\Phi(X)^{*}\Phi(X)$ are random variables taking value in $B(\mathcal{H},L^{2}),B(L^{2},\mathcal{H})$ and $B(\mathcal{H})$, respectively, and $\Phi(X)^{*}Y$ is a random variable taking value in $\mathcal{H}$.
@@
so that $\mathbb{E}\left[\lVert{\Phi(X)^{*}\Phi(X)}\rVert_{\mathcal{H},\mathcal{H}}\right]$ is finite and the expectation $\Sigma$ of $\Phi(X)^{*}\Phi(X)$ exists as Bochner integral in $\mathcal{B}(\mathcal{H})$. By a similar argument, $\widecheck{X}\convolution{X}$ is bounded in $L^{1}$ and its expectation exists as Bochner integral in $L^{1}$. By (12), $\Phi(X)^{*}Y=\Phi(X)^{*}\Phi(X)w_{*}+\Phi(X)^{*}\epsilon$. Since $=\Phi(X)^{*}\Phi(X)$ is bounded, so is $\Phi(X)^{*}\Phi(X)w_{*}$ and $\mathbb{E}\left[\lVert{\Phi(X)^{*}\Phi(X)w_{*}}\rVert_{\mathcal{H}}\right]$ is finite. By (14) and $\lVert{\Phi(X)^{*}}\rVert_{L^{2},\mathcal{H}}=\lVert{\Phi(X)}\rVert_{\mathcal{H},L^{2}}\leq D_{X}$, then
@@
Taking into account (48),
@@
then the random variable $\lvert{\mathcal{F}X}\rvert^{2}$, taking value in $\ell^{\infty}$, is bounded, so that it has finite expectation. Fix $\xi\in\widehat{G}_{*}$, by taking the expectation of (52), we get (57). ∎
@@
By using the operator $S_{n}$ the minimisation problem (30) reads as
@@
where $\mathbf{Y}=(Y_{1},\ldots,Y_{n})$. Hence, a standard result gives (70). Fix $\xi\in\widehat{G}_{*}$, by (68) and (69) then
@@
For $\xi\in\widehat{G}_{*}$, Eq. (71) is now consequence of (22). If $\xi\notin\widehat{G}_{*}$, Eq. (71) follows observing that $(\mathcal{F}j(w))_{\xi}=0$ for all $w\in\mathcal{H}$ by definition of $\mathcal{H}$. ∎
@@ A.3. Decomposition error
Eq. (A.5) is now consequence of the fact that $\lVert{\Sigma(\Sigma+\lambda\operatorname{I})^{-1}}\rVert^{\frac{1}{2}}_{\mathcal{H},\mathcal{H}}\leq 1$. The proof of (74) is similar by replacing the bound $\lVert{\Sigma(\Sigma+\lambda\operatorname{I})^{-1}}\rVert_{\mathcal{H},\mathcal{H}}^{\frac{1}{2}}\leq 1$ with $\lVert{(\Sigma+\lambda\operatorname{I})^{-1}}\rVert^{\frac{1}{2}}_{\mathcal{H},\mathcal{H}}\leq 1/\sqrt{\lambda}$. ∎
@@
Since both $\Sigma$ and $\Phi(X)^{*}\Phi(X)$ are trace class operators, so is $W$. Moreover, by (56) $\mathbb{E}\left[W\right]=0$ and, since $W\leq(\Sigma+\lambda\operatorname{I})^{-1}\Sigma$, then
@@
where the last inequality is a consequence of (55). Clearly
@@
then Eq. (91) with $t=1/2$ gives that
@@
since $\ln\ln y\geq 0$. This means that (84) holds true provided that
@@
which always holds true since, by (82)
@@
which by (14) satisfies $\mathbb{E}\left[Z\right]=0$. Moreover,
@@
where $M=M_{\epsilon}\kappa/\sqrt{\lambda}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}\operatorname{Tr}{\left((\Sigma+\lambda\operatorname{I})^{-1}\Sigma\right)}$. For any $i=1,\ldots,n$ set $Z_{i}=(\Sigma+\lambda\operatorname{I})^{-\frac{1}{2}}\Phi(X_{i})^{*}\epsilon_{i}$, then $Z_{1},\ldots,Z_{n}$ is a i.i.d. family of random variables distributed as $Z$ and, by (63),
@@
where $C$ is such that $\sigma_{\ell}\leq C^{-1}\ell^{-b}$. This shows (36). ∎
@@
Fix $\tau>1$ and $n\geq 3$. Assume that $\lambda>0$ satisfies (82) with $\delta=e^{-\tau}$, then, by Lemma A.6 and Lemma A.7, with probability at least $1-e^{-\tau}$,
@@
We pluggin (88) and (89) in (A.5) taking into account (86), so that with probability greater than $1-3e^{-\tau}$
@@
Assume that $(r,b)\neq(0,+\infty)$. Set $\lambda=\lambda_{n}$ as in (37), then taking into account
@@
with probability greater than $1-3e^{-\tau}$ provided that $n$ is large enough so that the right inequality in (82) holds true. This means that
@@
Let $n_{0}=n_{0}(\tau)\geq 3$ be the smallest integer such that the above inequality holds true, then (82) holds true forn any $n\geq n_{0}$ and this shows bound (38) if $(r,b)\neq(0,+\infty)$.
@@
so that (38) is clear by suitable definition of $n_{0}$. The proof of (39) is similar by using (74) instead of (A.5). ∎
@@ A.4. Technical results
For all $x\in L^{1}$ with $\lVert{x}\rVert_{1}\leq 1$, Young inequality (4) gives

### v3

@@ A.2. Probabilistic tools
Associated to the feature map $\Phi$, we introduce some useful random variables, which play a central role in the proofs. By eq. 52 the map $\Phi$ is continuous from $L^{1}$ into $\mathcal{B}(\mathcal{H},L^{2})$, then $\Phi(X),\Phi(X)^{*}$ and $\Phi(X)^{*}\Phi(X)$ are random variables taking value in $B(\mathcal{H},L^{2}),B(L^{2},\mathcal{H})$ and $B(\mathcal{H})$, respectively, and $\Phi(X)^{*}Y$ is a random variable taking value in $\mathcal{H}$.
@@
so that $\mathbb{E}\left[\lVert{\Phi(X)^{*}\Phi(X)}\rVert_{\mathcal{H},\mathcal{H}}\right]$ is finite and the expectation $\Sigma$ of $\Phi(X)^{*}\Phi(X)$ exists as Bochner integral in $\mathcal{B}(\mathcal{H})$. By a similar argument, $\widecheck{X}\convolution{X}$ is bounded in $L^{1}$ and its expectation exists as Bochner integral in $L^{1}$. By eq. 12, $\Phi(X)^{*}Y=\Phi(X)^{*}\Phi(X)w_{*}+\Phi(X)^{*}\epsilon$. Since $=\Phi(X)^{*}\Phi(X)$ is bounded, so is $\Phi(X)^{*}\Phi(X)w_{*}$ and $\mathbb{E}\left[\lVert{\Phi(X)^{*}\Phi(X)w_{*}}\rVert_{\mathcal{H}}\right]$ is finite. By eq. 13 and $\lVert{\Phi(X)^{*}}\rVert_{L^{2},\mathcal{H}}=\lVert{\Phi(X)}\rVert_{\mathcal{H},L^{2}}\leq D_{X}$, then
@@
Taking into account eq. 47,
@@
then the random variable $\lvert{\mathcal{F}X}\rvert^{2}$, taking value in $\ell^{\infty}$, is bounded, so that it has finite expectation. Fix $\xi\in\widehat{G}_{*}$, by taking the expectation of eq. 51, we get eq. 56. ∎
@@
By using the operator $S_{n}$ the minimisation problem in eq. 29 reads as
@@
where $\mathbf{Y}=(Y_{1},\ldots,Y_{n})$. Hence, a standard result gives eq. 69. Fix $\xi\in\widehat{G}_{*}$, by eqs. (67) and (68) then
@@
For $\xi\in\widehat{G}_{*}$, eq. 70 is now consequence of eq. 21. If $\xi\notin\widehat{G}_{*}$, eq. 70 follows observing that $(\mathcal{F}j(w))_{\xi}=0$ for all $w\in\mathcal{H}$ by definition of $\mathcal{H}$. ∎
@@ A.3. Decomposition error
Eq. (A.5) is now consequence of the fact that $\lVert{\Sigma(\Sigma+\lambda\operatorname{I})^{-1}}\rVert^{\frac{1}{2}}_{\mathcal{H},\mathcal{H}}\leq 1$. The proof of eq. 73 is similar by replacing the bound $\lVert{\Sigma(\Sigma+\lambda\operatorname{I})^{-1}}\rVert_{\mathcal{H},\mathcal{H}}^{\frac{1}{2}}\leq 1$ with $\lVert{(\Sigma+\lambda\operatorname{I})^{-1}}\rVert^{\frac{1}{2}}_{\mathcal{H},\mathcal{H}}\leq 1/\sqrt{\lambda}$. ∎
@@
Since both $\Sigma$ and $\Phi(X)^{*}\Phi(X)$ are trace class operators, so is $W$. Moreover, by eq. 55 $\mathbb{E}\left[W\right]=0$ and, since $W\leq(\Sigma+\lambda\operatorname{I})^{-1}\Sigma$, then
@@
where the last inequality is a consequence of eq. 54. Clearly
@@
then eq. 90 with $t=1/2$ gives that
@@
since $\ln\ln y\geq 0$. This means that eq. 83 holds true provided that
@@
which always holds true since, by eq. 81
@@
which by eq. 13 satisfies $\mathbb{E}\left[Z\right]=0$. Moreover,
@@
where $M=M_{\epsilon}\kappa/\sqrt{\lambda}$ and $\sigma^{2}=\sigma_{\epsilon}^{2}\operatorname{Tr}{\left((\Sigma+\lambda\operatorname{I})^{-1}\Sigma\right)}$. For any $i=1,\ldots,n$ set $Z_{i}=(\Sigma+\lambda\operatorname{I})^{-\frac{1}{2}}\Phi(X_{i})^{*}\epsilon_{i}$, then $Z_{1},\ldots,Z_{n}$ is a i.i.d. family of random variables distributed as $Z$ and, by eq. 62,
@@
where $C$ is such that $\sigma_{\ell}\leq C^{-1}\ell^{-b}$. This shows eq. 35. ∎
@@
Fix $\tau>1$ and $n\geq 3$. Assume that $\lambda>0$ satisfies eq. 81 with $\delta=e^{-\tau}$, then, by Lemma A.6 and Lemma A.7, with probability at least $1-e^{-\tau}$,
@@
We pluggin eqs. (87) and (88) in A.5 taking into account eq. 85, so that with probability greater than $1-3e^{-\tau}$
@@
Assume that $(r,b)\neq(0,+\infty)$. Set $\lambda=\lambda_{n}$ as in eq. 36, then taking into account
@@
with probability greater than $1-3e^{-\tau}$ provided that $n$ is large enough so that the right inequality in eq. 81 holds true. This means that
@@
Let $n_{0}=n_{0}(\tau)\geq 3$ be the smallest integer such that the above inequality holds true, then eq. 81 holds true for any $n\geq n_{0}$ and this shows bound (37) if $(r,b)\neq(0,+\infty)$.
@@
so that eq. 37 is clear by suitable definition of $n_{0}$. The proof of eq. 38 is similar by using eq. 73 instead of A.5. ∎
@@ A.4. Technical results
For all $x\in L^{1}$ with $\lVert{x}\rVert_{1}\leq 1$, Young inequality in eq. 4 gives